0 0
Add a comment Improve this question Transcribed image text
Answer #1

5) Velocity tripled so K_{f} = 9K_{i}

Total work done = K_{f} - K_{i}= 8K_{i} , Option B

6) We dont apply any force in the direction of displacement, only perpendicular to it. So no work done. Option B.

7) Now work is done against gravity = m*g*h = 14*1.7*9.8 = 233.24J . Option A is close